OpenStudy (anonymous):

Sofia is buying new skis. She finds that the average price of skis is $500 but the actual price could differ from the average as much as $250. Write and solve an absolute value inequality to describe this situation.

OpenStudy (anonymous):

\[|x - 500| \le 250\]\[-250 \le x -500 \le 250\] the difference between the actual price and the average is less than or equal to 250